OpenStudy (anonymous):

What is the vertex of the graph of y = 5(x + 4)2 + 3? A.(–4, 3) B.(4, 3) C.(3, 4) D.(3, –4)

OpenStudy (raden):

y = 5(x + 4)2 + 3 y = 5(x -(- 4))^2 + 3 now, looks obviously the vertec is (-4,3)